50 gram of water at 0°C is mixed with an equal mass of water at 80°C. Calculate the resultant increase in entropy


1
Expert's answer
2021-04-01T18:33:21-0400

Given,

mass of the water = 50gm

Temperature of the water "T_1=0^\circ C"

mass of the second water =50gm

Temperature of the water "(T_2)=80^\circ C"

Let the final temperature of the water be t.

Now, applying the conservation of energy,

Heat gain = Heat loss

"50s(t-0)=50s(80-t)"

"t=40^\circ C"

Change in entropy of "0^\circ C" water = "\frac{\Delta Q}{t}=\frac{50\times 40}{40+273}=6.39J/k"

Similarly change in entropy of "80^\circ C" water,

"=-\frac{50\times 40}{40+273}=-6.39J/k"

Hence the total change in entropy=0J/k